OpenStudy (anonymous):

The equation of the tangent line to the curve x^2 + y^2 = 169 at the point (5, -12) is...

OpenStudy (anonymous):

First, take the derivative. Then, evaluate the derivative at (5,-12) Then, put the slope from the derivative and the point (5,-12) into the point slope form for a line.

OpenStudy (anonymous):

I got the slope as -5/6?

OpenStudy (anonymous):

I'm thinking of something more like -5/12. Use implicit differentiation.

OpenStudy (anonymous):

oh ok. I have that now. So for the next step I pluf in 5 and 12 for x and y?

OpenStudy (anonymous):

jk. already did that

OpenStudy (anonymous):

sorry, I'm stuck again

OpenStudy (anonymous):

With?

OpenStudy (anonymous):

"Then, put the slope from the derivative and the point (5,-12) into the point slope form for a line."

OpenStudy (anonymous):

Point slope form for a line, with slope m, , going through a point:\[(x_1,y_1)\] is \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

oh,ok. but I still can't seem to get an answer thats one of my options. I got 12y - 5x = -37

OpenStudy (anonymous):

Try again, I'm getting something different.

OpenStudy (anonymous):

oh ok... I got 12y + 5x = -169

OpenStudy (anonymous):

-5x

OpenStudy (anonymous):

This is much closer to what I found, though I question your alteration of signs.

OpenStudy (anonymous):

well my final answer was 5x - 12y = 169. Is that still wrong?

OpenStudy (anonymous):

I end up with 12y+5x=-119, without checking my work, writing it down, etc...so it may be wrong.

OpenStudy (anonymous):

looking over this conversation, you used slope as -5/12 but I used 5/12 b/c isnt y' = -x/y? and -5/-12 = positive?

OpenStudy (anonymous):

Absolutely. Yet again, I left out a negative in my original derivative. I really should just get some sleep.

OpenStudy (anonymous):

ok, well at least I know the process now. I can check my work. Thank you for your patience!

OpenStudy (anonymous):

C'est rien.